OpenStudy (anonymous):

In parallelogramDEFG,DH=y+3,HF=2x+1,GH=2y–1,andHE=x+4.Find the values of x and y

OpenStudy (anonymous):

Please help me,

OpenStudy (anonymous):

x=35; y=20

OpenStudy (anonymous):

1) DH = HF x + 5 = 2y x = 2y - 5 2) GH = GE 3x - 1 = 5y + 4 3x = 5y + 5 x = (5y + 5)/3 2y - 5 = (5y + 5)/3 6y - 15 = 5y + 5 y - 20 = 0 y = 20 x = 2y - 5 x = 2(20) - 5 x = 35 check: 3x - 1 = 5y + 4 3(35) - 1 = 5(20) + 4 105 - 1 = 100 + 4 104 = 104

OpenStudy (anonymous):

can i have medal and fan

OpenStudy (anonymous):

There isn't any answer choice of 35 and. 20 these are the ones x=14, y=11. X=3 and y=4. X=7 and y=4. X=1 and y=11

OpenStudy (anonymous):

If you help me with this I will give you fan and medal

OpenStudy (anonymous):

Hello!

RhondaSommer (rhondasommer):

i can try to help

RhondaSommer (rhondasommer):

are you sure the parellelgam is DEFG ? deosnt make sense if there is an F and H

OpenStudy (triciaal):